# Novel Challenges in Tracking Self-Interacting Dark Matter Subhalos

### v1

@@ 1 Introduction
In the standard Lambda Cold Dark Matter ($\Lambda$CDM) cosmological model, dark matter (DM) halos originate from primordial density fluctuations. Large structures form hierarchically through the merging of numerous small halos. During this merging process, halos that are accreted into a larger halo still survive as self-bounded “subhalos” of their hosts [1, 2, 3, 4, 5]. Predictions for DM substructure are crucial to test our current understanding of structure formation on small scales and to interpret observations of satellite galaxies, stellar streams, and strong-lensing systems. Importantly, the distribution and evolution of substructure are sensitive to the particle nature of DM, since physics beyond gravity can change subhalo abundances and density profiles [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23]. Subhalos are therefore critical for testing DM on sub-galactic scales and addressing the diverse distribution of DM in galaxies, which is currently a prominent challenge to the $\Lambda$CDM framework [24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44].
@@
Several recent studies have developed different software tools and numerical techniques to address these problems. Of particular relevance for this paper are friends-of-friends algorithms that group particles belonging to distinct phase-space overdensities, and particle-tracking algorithms that follow some subset of a subhalo’s particles forward in time, e.g., after infall into a larger host [59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 58, 55]. Multiple particle-tracking-based halo finders have been developed based on a similar underlying algorithm that first associates particles with subhalos at a certain time, using a phase-space friends-of-friends algorithm, and then tracks particles forward to determine descendant subhalos in later snapshots. Compared to algorithms that solely rely on phase-space information, particle-tracking approaches recover more subhalos and are able to follow subhalos longer, although the exact performance varies between different simulations and halo finders [64, 67, 66, 55, 58, 71]. However, even though the particle-tracking technique generally improves the accuracy of DM substructure predictions, its fidelity should be revisited in scenarios involving DM physics beyond gravity. In these scenarios, the tidal evolution of subhalos can differ from CDM expectations, and all particle-tracking-based algorithms involve certain numerical choices and assumptions that are not always physically motivated.
@@
[…]
@@ 2.2 Subhalo Finders and Catalogs
We compare the Symfind results to the post-processed catalogs from the Rockstar subhalo finder [65] and the Consistent-Trees merger tree code pipeline [108]. The catalog post-processing is also performed by Symfind, as described in [58], to enable one-to-one subhalo matching across different catalogs. Furthermore, for each of the CDM subhalos, we identify its SIDM counterpart to allow for one-to-one comparisons between their evolution and trajectories. Following [18, 19], we take the same definition of the virial mass as $\Delta_{\rm vir}\approx 99$ times the critical density of the Universe at $z=0$ for the cosmological parameters adopted in the simulations [109]. We also adopt the same definition for the peak mass $M_{\rm peak}$ as in [58], i.e., the maximum mass value before a subhalo ever became a subhalo of a larger system. This prevents numerical spikes in mass found in subhalo finders like Rockstar or Subfind from biasing $M_{\rm peak}$. For simplicity, we will use “RCT” to denote the results from the Rockstar+Consistent-Trees pipeline for the rest of the paper.
@@ 3.1 Core-expansion SIDM Subhalos
In the bottom panel of figure 2, we show the evolution of the distance from the core particles to the center of the subhalo. For the CDM subhalo (blue), there is a very mild overall trend that causes the core particles to diffuse outward over time. This is likely caused by numerical two-body relaxation in the halo’s center [e.g. 111, 112]. In contrast, the SIDM counterpart’s core particles migrate to much larger distances over time due to a combination of self-interaction and tidal stripping. After tidal stripping, some can be found hundreds of kpc away from the center of the subhalo.
@@ 4 Subhalo Mass Functions
We have also checked that increasing the number of core particles from $N_{\rm core}=32$ to $128$ in Symfind does not improve the subhalo population recovery in the high-$M_{\rm peak}$ regime for the GroupSIDM model. On the contrary, increasing the core particles leads to fewer recovered subhalos with $N_{\rm peak}>10^{3.5}$. This is likely because Symfind’s method for determining subhalo position is based on finding the density peak that is associated with the largest fraction of core particles. Most cases of disruption are due to the true remnant containing relatively few core particles relative to the extended density peak represented by the tidal tails and extended phase-mixed mass component, and increasing core particle count is unlikely to help. Once the core particle count is large enough to prevent shot-count effects, increasing the particle count likely only dilutes the core particle pool with less bound particles.
@@ 6 Conclusions
[…]
@@
The loss of core particles has a more significant implication on using “orphan” models in SIDM simulations, which rely on one or more most-bound particles represent subhalos [60, 61, 62, 63, 69, 70, 72, 131, 73], as such assumptions may not hold in SIDM. We note that it will be essential to develop similar tests for hydrodynamic simulations in both CDM and SIDM. In particular, non-gravitational physics (e.g., supernova feedback) can presumably also lead to the unbinding of core particles in these scenarios, which we will leave for future work.
@@ Appendix B Evolution of Core Particles in Idealized Simulations
[…]

### v2

@@ 1 Introduction
[…]
@@
Several recent studies have developed different software tools and numerical techniques to address these problems. Of particular relevance for this paper are friends-of-friends algorithms that group particles belonging to distinct phase-space overdensities, and particle-tracking algorithms that follow some subset of a subhalo’s particles forward in time, e.g., after infall into a larger host [59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 58, 55, 74]. Multiple particle-tracking-based halo finders have been developed based on a similar underlying algorithm that first associates particles with subhalos at a certain time, using a phase-space friends-of-friends algorithm, and then tracks particles forward to determine descendant subhalos in later snapshots. Compared to algorithms that solely rely on phase-space information, particle-tracking approaches recover more subhalos and are able to follow subhalos longer, although the exact performance varies between different simulations and halo finders [64, 67, 66, 55, 58, 71]. However, even though the particle-tracking technique generally improves the accuracy of DM substructure predictions, its fidelity should be revisited in scenarios involving DM physics beyond gravity. In these scenarios, the tidal evolution of subhalos can differ from CDM expectations, and all particle-tracking-based algorithms involve certain numerical choices and assumptions that are not always physically motivated.
@@
Thus, SIDM physics can influence the relative performance of subhalo finders in a model-dependent fashion, which has important implications for SIDM analyses. For example, particle-tracking-based methods can underpredict the abundance of bright satellite galaxies by missing stripped, core-expansion subhalos, but they generally identify more core-collapse subhalos than phase-space methods, which is key for interpreting ultra-faint dwarfs and lensing perturbers. Given these complex tradeoffs, we ultimately recommend using a combination of both algorithms to track SIDM subhalos and construct the most robust SIDM subhalo catalogs. In our subhalo mass function tests for two host halos with masses $10^{12}~{\rm M_{\odot}}$ and $10^{13}~{\rm M_{\odot}}$, the combined catalogs contain $2\textup{--}15\%$ more subhalos than either method alone when considering the overall population, and $10\textup{--}30\%$ more subhalos in the inner region of the host halo. For this reason, the subhalo catalogs in SIDM Concerto [23] are constructed using both methods.
@@ 2.2 Subhalo Finders and Catalogs
We compare the Symfind results to the post-processed catalogs from the Rockstar subhalo finder [65] and the Consistent-Trees merger tree code pipeline [109]. The catalog post-processing is also performed by Symfind, as described in Appendix A of [58]. We note that this post-processing step is key to enabling one-to-one subhalo matching with Rockstar catalogs. In particular, this step takes results from Rockstar-plus-Consistent-Trees as the initial input and re-processes the merger tree by removing false subhalos, unphysical mass accretion histories, and incorrect subhalo property measurements caused by Rockstar losing track of halos as they disrupt. The current version of Symfind focuses on all (sub)halos that ever orbit within $R_{\rm vir}$ of the host system. Thus, Symfind re-indexes (sub)halos based on their largest pre-infall mass, using information from the post-processed Rockstar catalog, thereby aligning Rockstar and Symfind catalogs. As we will show later, unique halo IDs from the Rockstar catalog assigned at simulation time are converted into Symfind indices for easier access (e.g., a subhalo with index $10$ at snapshot $z=2$ is the primary progenitor of the subhalo with index $10$ at snapshot $z=0$).
On the other hand, the Symfind algorithm only begins operating at the first snapshot in which a (sub)halo enters within $R_{\rm vir}$ of the host system; it then tracks the subhalo through to the end of the simulation. As we show in the next section, the Symfind catalogs contain information only after first infall, while adopting all information from Rockstar catalogs beforehand. Consequently, cross-comparisons between Rockstar and Symfind catalogs are one-to-one for the same subhalo (e.g., a subhalo with index $10$ in the post-processed Rockstar catalog corresponds to the subhalo with index $10$ in the Symfind catalog). We therefore use subhalo indices to combine Symfind and Rockstar catalogs without double-counting any subhalos. In cases where both Symfind and Rockstar identify a given subhalo, we adopt the subhalo properties measured by Rockstar.
Furthermore, for each of the CDM subhalos, we identify its SIDM counterpart to allow for one-to-one comparisons between their evolution and pre-infall trajectories. Following [18, 19], we take the same definition of the virial mass as $\Delta_{\rm vir}\approx 99$ times the critical density of the Universe at $z=0$ for the cosmological parameters adopted in the simulations [110]. We also adopt the same definition for the peak mass $M_{\rm peak}$ as in [58], i.e., the maximum mass value before a subhalo ever became a subhalo of a larger system. This prevents numerical spikes in mass found in subhalo finders like Rockstar or Subfind from biasing $M_{\rm peak}$. For simplicity, we will use “RCT” to denote the results from the Rockstar+Consistent-Trees pipeline for the rest of the paper.
@@ 3.1 Core-expansion SIDM Subhalos
In the bottom panel of figure 2, we show the evolution of the distance from the core particles to the center of the subhalo. For the CDM subhalo (blue), there is a very mild overall trend that causes the core particles to diffuse outward over time. This is likely caused by numerical two-body relaxation in the halo’s center [e.g. 112, 113]. In contrast, the SIDM counterpart’s core particles migrate to much larger distances over time due to a combination of self-interaction and tidal stripping. After tidal stripping, some can be found hundreds of kpc away from the center of the subhalo.
@@ 4 Subhalo Mass Functions
We have also checked that increasing the number of core particles from $N_{\rm core}=32$ to $128$ in Symfind does not improve the subhalo population recovery in the high-$M_{\rm peak}$ regime for the GroupSIDM model. On the contrary, increasing the number of core particles leads to fewer recovered subhalos with peak particle counts $N_{\rm peak}>10^{3.5}$. This is likely because Symfind’s method for determining subhalo position is based on finding the density peak that is associated with the largest fraction of core particles. Most cases of disruption are due to the true remnant containing relatively few core particles relative to the extended density peak represented by the tidal tails and extended phase-mixed mass component, and increasing core particle count is unlikely to help. Once the core particle count is large enough to prevent shot-count effects, increasing the particle count likely only dilutes the core particle pool with less bound particles.
@@ 6 Conclusions
Since the performance of subhalo finders depends on the underlying SIDM model, and there is no clear preference for solely using Symfind or RCT, we recommend using a combination of both methods to construct the most robust SIDM subhalo catalogs. In our subhalo mass function tests, we find that the combined catalogs contain $2\textup{--}15\%$ more subhalos than either method alone when considering the overall population, and $10\textup{--}30\%$ more subhalos in the inner region ($r<R_{\rm vir}/4$) of the host system.
@@
The loss of core particles has a more significant implication on using “orphan” models in SIDM simulations, which rely on one or more most-bound particles represent subhalos [60, 61, 62, 63, 69, 70, 72, 132, 73], as such assumptions may not hold in SIDM. We note that it will be essential to develop similar tests for hydrodynamic simulations in both CDM and SIDM. For example, a recent study shows that collisional particles (such as gas particles in hydrodynamic simulations) may not be reliable tracers for tracking subhalos [74]. Furthermore, non-gravitational physics (e.g., supernova feedback) can presumably also lead to the unbinding of core particles in these scenarios, which we will leave for future work.
@@ Appendix B Evolution of Core Particles in Idealized Simulations
In figure 11, we show the distance of core particles from the halo center in idealized CDM (blue) and SIDM (magenta) simulations. In the CDM halo, nearly all core particles remain within $1~{\rm kpc}$ of the center at all times. In contrast, core particles in the SIDM halo are dispersed over a much wider range. Most of them can migrate to $20~{\rm kpc}$ and a few of them even reach $80\textup{--}100~{\rm kpc}$. Since tidal stripping is absent in the isolated simulations, the difference in the distribution of core particles between SIDM and CDM halos is entirely due to DM self-interactions. We have also verified that the core size of the simulated SIDM halo is $20~{\rm kpc}$, consistent with the migration range of the core particles.